OpenStudy (anonymous):

help...Write the equation of the line that is parallel to the line y = 2x + 2 and passes through the point (5, 3).

OpenStudy (jonnyvonny):

For a line to be parallel to another, what 2 conditions must be met?

OpenStudy (anonymous):

The answer choices are... y = 2x -7 y = -1/2x - 2 y = 2x - 2 y = - 1/2x -7

OpenStudy (jonnyvonny):

Since you're given the answers, all you really must to is know that the slopes MUST be the same, that is 2, in this case. From there, you can just plug in x=5 for the choices that have the slope equaling 2, and see if the result is 3. Do you wish tot know how to do this problem if the answers were not given?

OpenStudy (anonymous):

not really but you can if you want to

OpenStudy (anonymous):

so what would the answer be

OpenStudy (anonymous):



OpenStudy (anonymous):

so A

OpenStudy (jonnyvonny):

Yes.